OpenStudy (kamille):

How many µl are there in 0,75ml? a) 750 b) 75 c) 7.5 d)7500 Can anyone explain?

OpenStudy (phi):

Do you know what the µ stands for in µl ?

OpenStudy (kamille):

yeah, it is micro, which means \[1000^{-2}*1000=1000\] ml?

OpenStudy (phi):

ok, µ means micro, and micro means 1 millionth, 1 µl is \[ \frac{ 1 \text{ liter}}{1000000} \] it is tough to count the zeros, so people write it as \[ \frac{ 1 \text{ liter}}{10^6} \] or, "moving" the 10^6 up top (and changing the 6 to -6) 1 µl= \(10^{-6} \) liters

OpenStudy (phi):

do you know m in ml stands for milli (one thousandth) \[ 1 ml= \frac{1 \text{ liter }}{1000} = 10^{-3} \text{ liters} \]

OpenStudy (kamille):

oh,yeah. So now I need to convert 0,75ml into µl?

OpenStudy (phi):

to solve, multiply 0,75ml by the number of µl in 1 ml (1000 or 10^3)

OpenStudy (kamille):

750, right!?

OpenStudy (phi):

yes

OpenStudy (kamille):

THANKS! are you good at biochemistry, maybe? i desperately need help and none can help me with one question...